In a right triangle, one acute angle is 54° more than twice the other. Find each acute angle.
Answer by orca(409)

Let x represent the smaller angle.
Then the larger one is 2x + 54.
As the sum of the measures of the two angles is 90, we have
x + 2x + 54 = 90
Solving for x, we have
3x + 54 = 90
3x = 90-54
3x = 36
x = 12
So the smaller angle is 12 degrees
The larger one is 2x + 54 = 2x12 + 54 = 78 degrees